OpenStudy (anonymous):

help medal quick In what direction is the line containing the points (-10, -6) and (-1, 8) going?

OpenStudy (steve816):

Find the slope :) If the slope is positive, then it is going upwards. If the slope is negative, it is going downwards.

OpenStudy (shaik0124):

slope=8+6/-1+10 =14/9